What is the equation of the line described below written in slope-intercept form?
the line passing through point (0, 4), perpendicular to the line whose equation is y =1/-4x

y = -4x
y = -4x + 4
y = x + 4
y = 4x + 4

Answer by stanbon(75887)

What is the equation of the line described below written in slope-intercept form?
the line passing through point (0, 4), perpendicular to the line whose equation is y =1/-4x
-----
The given line has slope = -1/4
------
The perpendicular line must have slope = 4
-------
Form:: y = mx + b
Since f(0) = 4
---------------
Equation:
y = 4x + 4
------------------
Cheers,
Stan H.
===============
y = -4x
y = -4x + 4
y = x + 4
y = 4x + 4